OpenStudy (laddiusmaximus):

calculate the derivative of y in respect to x x^(3)y+3xy^(3)=x+y

OpenStudy (anonymous):

guess my answer was not sufficient. do you know how to use implicit diff?

OpenStudy (laddiusmaximus):

no. the teacher covered it in class today...sort of.

OpenStudy (anonymous):

\[x^ 3y+3xy^3=x+y\] think of \[x^3f(x)+3xf^2(x)=x+f(x)\]

OpenStudy (anonymous):

you need to use the product rule to find the derivative of \[x^3f(x)\] it is \[3x^2f(x)+x^3f'(x)\] then replacing f(x) by y you get \[3x^2y+x^3y'\]

OpenStudy (anonymous):

for \[3xy^3\] you need both the product rule and the chain rule

OpenStudy (anonymous):

the derivative of \[3xf^3(x)\] is \[3f^3(x)+9xf^2(x)f'(x)\] or the derivative of \[3xy^3\] is \[3y+9xy^2y'\]

OpenStudy (anonymous):

In this case, implicit differentiation has to be used. So, you are finding derivative of y respect to x x^(3)y+3xy^(3)=x+y First, Find the derivative of x^3 y by using product rule Derivative of x^3 y is 3x^2 y + dyx^3 . Then, find the derivative of 3xy^3 with the same rule Derivative of 3xy^3 is 3y^3 + 3y^2 (dy) (3x) = 3y^3 + 9xy^2 (dy) After that, find the derivative from the other side of equation and you get, 1 + dy. Let's write down everything back, 3x^2 y + dyx^3 + 3y^3 + 9xy^2 (dy) = 1 + dy Send all of the dy to one side, dyx^3 + 9xy^2 (dy) - dy = 1- 3x^2 y - 3y^3 Factor out the dy 's, dy ( x^3 + 9xy^2 - 1) = 1 - 3x^2 y - 3y^3 Then, divide both side with ( x^3 + 9xy^2 - 1) So you have dy = [ 1- 3x^2 y - 3y^3 ] / [x^3 + 9xy^2 - 1] Hope, above helps.. srry if it's too long ...heh